# RUMOR: Reinforcement learning for Understanding a Model of the Real World for Navigation in Dynamic Environments

### v1

@@ 6.2 Simulation experiments
RUMOR underperforms NR-RUMOR as expected, as kinematic and dynamic restrictions limits the capacity of maneuvering and sudden reactions. Both success and time metrics are worse in dense scenarios, whereas, even though the success rate is slightly higher, the navigation time is clearly higher in low occupancy scenarios. Nonetheless, their performance is comparable despite the restrictions. Furthermore, it displays a better performance, or at least comparable, than the rest of the planners, showing to be the second best overall. The difference in the behavior is seen in the velocity and path length graphs. While RUMOR tries to reach the maximum linear velocity as the rest of the planners, the acceleration limits prevents it from achieving it as much time as the rest. In addition, it has to keep a lower velocity in dangerous zones, as it must stop gradually. The path length and navigation time plots show that RUMOR achieve high success rates with longer paths, meaning that it must respect the restrictions to navigate, it follows safer paths where no sudden velocity changes are needed.
@@ 6.3 Real-world experiments
Real-world experiments were conducted to validate and evaluate RUMOR performance in real conditions. The whole system was integrated in a Turtlebot 2 platform with a NUC with Intel Core i5-6260U CPU and 8 GB of RAM. The sensor used is a 180º Hokuyo 2D-LIDAR. Due to the real-world design of the system, the same network weights and ROS nodes used in simulation were used in the ground robot, adapting the obstacle tracker node to detect people and other objects and using AMCL for localization. Two different types of real-world experiments were designed. Those experiments may be seen in the supplementary video.
@@
Finally, a setup different from a lab scenario was tried, in settings including a corridor with some rooms and intersections, obstacles with different shapes and people with heterogeneous behaviors, as in real life, to prove that the system works in a completely unseen environment that has not been specifically used in simulation. This may be seen in Figure 10. A global planner as the one introduced in [45] was used to help the robot navigate through rooms and convex obstacles, while letting RUMOR compute the motion commands by itself. The results show that the robot is able to navigate very smoothly through the dynamic obstacles that have behaviors different from those seen in training, avoids static obstacles that had not been seen before, and works with the global planner to be able to reach goals that are very far from the initial position. The use of DOVS, instead of being a complete end-to-end planner, limits out-of-distribution observations, as apparently different scenarios could result in similar DOVS representations.
@@ 7 Conclusion
This work presents a novel motion planner for dynamic environments. It uses the DOVS model to extract the dynamism of the environment, abstracting from specific sensor data directly used by end-to-end approaches, and DRL to select motion commands that efficiently leads the robot to a goal while avoiding collisions with moving obstacles. Instead of directly use the raw obstacle information, which may lead to out-of-distribution observations in scenarios different from the ones seen in training, the system uses DOVS, which encodes it in similar terms regardless the scenario. In addition, we propose a training framework that puts the agent in real-world setting, an advanced DRL algorithm that uses memory to mitigate partial observability issues and a novel action space that intrinsically considers differential-drive kinodynamics, in order to mitigate sim2real transition. The proposed system is tested in random scenarios with absolute perception (ground truth from the simulator) and partial observability conditions, outperforming existing methods and showing the benefits of combining a model-based approach with a DRL controller. The algorithm is also tested in a ground robot in dense, dynamic and heterogeneous scenarios. Future work could include increasing the robustness of the method in scenarios where the obstacle motion is very irregular and unnatural, extending the model to collaborative collision avoidance with multi-robot navigation or adapting the model for 3-D navigation and UAVs.
@@
This work was partially supported by the Spanish projects MCIN/AEI/10.13039/501100011033/FEDER-UE, and Aragon Government_FSE-T45_23R.

### v2

@@ 5.2 Simulation experiments
RUMOR underperforms NR-RUMOR as expected, as kinematic and dynamic restrictions limits the capacity of maneuvering and sudden reactions. Nonetheless, it displays a better performance, or at least comparable, than the rest of the planners. The difference in the behavior is seen in the velocity and path length graphs. While RUMOR tries to reach the maximum linear velocity as the rest of the planners, the acceleration limits prevents it from achieving it as much time as the rest. In addition, it has to keep a lower velocity in dangerous zones, as it must stop gradually. The path length and navigation time plots show that RUMOR achieve high success rates with longer paths, meaning that it must respect the restrictions to navigate. It follows safer paths where no sudden velocity changes are needed.
TEB and DWA performance is the same in both perception settings, as their input, which is the sensor scan and not the individual agent information, is the same in both cases. TEB suffers from the same issues as the other model-based planners. DWA performance is comparable or better than other methods when they suffer from perception errors. It is similar to the restricted RUMOR in those cases. Nevertheless, in absolute perception settings, RUMOR outperforms DWA in both success rate and time to reach the goal with statistical significance, even considering Remark 3. Thus, it seems like using more accurate perception algorithms (e.g. deep learning) would increase the difference between both methods.
Two key components of the system explain NR-RUMOR results. First, the input of RUMOR network is the DOVS model of the environment. With 12 surrounding dynamic obstacles, there are unlimited number of possible scenario configurations, and the agent can not be trained to have a robust performance in all of them. On the contrary, DOVS represents the environment in similar terms no matter the scenario, leading to a deeper understanding of the velocity space. When the robot learns to interpret DOVS, it knows how to avoid collisions regardless the amount, positions or velocities of the obstacles. Second, the memory layer of RUMOR network allows it to keep track of previous observations. This is important to overcome partial observability challenges and those derived from sudden changes in the obstacles behavior. Overall, considering every combination in the perception and occupancy settings, NR-RUMOR consistently shows the best performance when compared with the other methods.
@@ 5.3 Real-world experiments
Hardware experiments were conducted to validate and evaluate RUMOR performance in real conditions. Two different types of real-world experiments were designed, which may be seen in the supplementary video.
@@
Finally, a setup different from a lab scenario was tried, in settings including a corridor with some rooms and intersections, obstacles with different shapes and people with heterogeneous behaviors, as in real life, to prove that the system works in a completely unseen environment that has not been specifically used in simulation. This may be seen in Figure 10. The results show that the robot is able to navigate very smoothly through the dynamic obstacles that have behaviors different from those seen in training, avoids static obstacles that had not been seen before, and works with the global planner to be able to reach goals that are very far from the initial position. The use of DOVS, instead of being a complete end-to-end planner, limits out-of-distribution observations, as apparently different scenarios could result in similar DOVS representations.
@@ 6 Conclusion
This work presents a novel motion planner for dynamic environments. It uses the DOVS model to extract the dynamism of the environment, abstracting from specific sensor data directly used by end-to-end approaches, and DRL to select motion commands that efficiently leads the robot to a goal while avoiding collisions with moving obstacles. Instead of directly use the raw obstacle information, which may lead to out-of-distribution observations in scenarios different from the ones seen in training, the system uses DOVS, which encodes it in similar terms regardless the scenario. In addition, we propose a training framework that puts the agent in real-world setting, an advanced DRL algorithm that uses memory to mitigate partial observability issues and a novel action space that intrinsically considers differential-drive kinodynamics, in order to mitigate sim2real transition. The proposed system is tested in random scenarios with absolute perception (ground truth from the simulator) and partial observability conditions, outperforming existing methods and showing the benefits of combining a model-based approach with a DRL controller. The algorithm is also tested in a ground robot in dense, dynamic and heterogeneous scenarios.
Future work could include increasing the robustness of the method in scenarios where the obstacle motion is very irregular and unnatural, extending the model to collaborative collision avoidance with multi-robot navigation or adapting the model for 3-D navigation and UAVs. In addition, we believe that studying the impact of the perception errors in both training and deployment separately would be of great interest; considering a deeper study of different perception algorithms, the effect in neural networks with different complexity, the inclusion of explicit uncertainty estimation and the evolution of the DOVS in presence of noise.
@@
This work was partially supported by MICIU/AEI/ 10.13039/501100011033 and ERDF/EU by grant PID2022-139615OB-I00 and grant PRE2020-094415, and Government of Aragón under grant DGA T45-23R.